Which of the following equations cannot be formed using the equation x = 7? (a) 2x + 1 = 15 (b) 7x – 1 = 50 (c) x – 3 = 4 (d) (x/7) – 1 = 0

Grade:12

1 Answers

Harshit Singh
3 years ago

Dear Student

Option (b) is correct here
Consider the equation 7x–1 = 50.
Transpose -1 to right hand side then it becomes 1.

7x = 50 + 1
7x = 51
x = 51/7

Thanks